OpenStudy (anonymous):

Can someone please explain to me cos, tan & sin? I'm doing advanced math and we're going to be doing these..most of my peers know a bit about this, and I don't want to fall behind. So an introduction on these would be great, and also maybe some problems which use these.. ?

OpenStudy (chaise):

Explain cos, tan & sin with respect to right angle triangles? Explain cos, tan & sin and their derivatives/integrals? Explain cos, tan & sin with respect to periodic functions? Explain cos, tan & sin with irregular triangles (non right angled) Anything else I am missing, if you've never encountered sin, cos or tan before in your mathematic life it would be best to start with right angle triangles. I can help with that, if that's what you're asking.

OpenStudy (anonymous):

Yeah, basically the basics. We're JUST ABOUT to learn about them now..I've heard about them, but don't know what they are exactly.

OpenStudy (anonymous):

So it'd be great if you start with the right angle trriangle cos/tan/sin :)

OpenStudy (chaise):

 Cos, tan and sin are simply ratios of the given lengths and their angles. First off you need to know how right angled triangles are labelled. The hypotenuse (h) in a right angled triangle is usually denoted with a h. The length opposite the angle in the right angle triangle (theta) is called opposite (o), and the odd one out (the length adjacent to the angle) is called the adjecent (a). Simple ratios combined the angle with their given lengths, and if the lengths are known, the angle can be determined. This will make more sense as you learn more about right angled triangles. \[\tan \theta=\frac{length of opposite side}{length of adjacent side}\] \[\sin \theta=\frac{length of opposite side}{length of hypotenuse}\] \[\cos \theta {length of adjacent side}{length of hyptenuse}\]